OpenStudy (anonymous):

Determine real roots: (3x^2+27)(x^2-16) i don't understand how to solve this, please help ! :)

OpenStudy (anonymous):

What values for x make the expression equal to zero?

OpenStudy (anonymous):

how do i figure that out ?

OpenStudy (anonymous):

typos...

OpenStudy (anonymous):

i know, but what i mean is that what method or technique is used here to figure it out

OpenStudy (anonymous):

when (x^2-16) = 0 or (3*x^2+27) =0

OpenStudy (anonymous):

that's it really...solve for x

OpenStudy (anonymous):

x^2=16 x=? x^2 =-9 x=?

OpenStudy (anonymous):

glad I could help